If I have a 97 average on my test which is 75% of my grade. The final is worth 25%, What do I need to make on the final to get a 91 in the class?

Answer by TimothyLamb(4379)

97(0.75) + f(0.25) = 91
72.75 + f(0.25) = 91
f(0.25) = 91 - 72.75
f(0.25) = 18.25
f = 73
---
must get 73% on final to get a class grade of 91%
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php